# Gen-C: Populating Virtual Worlds with Generative Crowds

### v3

@@ 4.2 Graph-Based Representation
Training a generative crowd model requires an expressive data structure that captures agent–environment interactions and actions over time. We adopt a graph representation that meets these criteria and introduce the crowd scenario graph, a time-expanded graph modeling dynamic crowd scenarios through temporally encoded actions and interactions.
Figure 1: Snapshot of a crowd scenario graph. Nodes encode an action, a location, and an agent ID with timestep (i_ti\_t).
As described in Section 4.1, during simulation we collect $Rec_{i}$ for each agent $i$. Iterating through $Rec_{i}$, we create a node $V_{i}^{t}=(i,A_{i}^{t},L_{i}^{t})$, where $t$ is the current timestep, $A_{i}^{t}\in\mathrm{A}ct$, and $L_{i}^{t}\in\mathrm{L}oc$; the combination of $A_{i}^{t}$ and $L_{i}^{t}$ defines an agent-environment interaction at $t$. We set the action sequence of $i$ using a “sequence edge” $E_{i}^{t}=(V_{i}^{t},V_{i}^{t-1})$. Next, agent-agent interactions are captured by connecting nodes from different agents that share an action at a given timestep. A “share edge” $E_{i,j}^{t}=(V_{i}^{t},V_{j}^{t})$ is created for each pair of agents $i$ and $j$ interacting at $t$. At this stage, nodes of each agent $i$ are sequentially connected into a chain. If the chain has no “shared edges,” it forms an individual subgraph $H_{k}$ corresponding to agent $i$. When at least one “shared edge” exists, the chains of multiple agents merge into a subgraph representing a group of agents. Thus, for each $CS$, we construct an undirected graph $G=(V,E)$, consisting of multiple subgraphs $\mathcal{H}=\{H_{1},H_{2},\dots\}$, where $V$ is the set of nodes and $E$ the set of edges. Finally, each subgraph $H_{k}\in\mathcal{H}$ is treated as an individual training sample $(S_{in},H_{k},\text{agents}_{k})$, where $H_{k}$ is the subgraph for group $k$, and $\text{agents}_{k}$ the agents in that group. Figure 1 presents a snapshot of a subgraph with five agents from a crowd scenario graph.
@@ 5.1 Data Preprocessing
Node Ordering. We enforce a canonical ordering of nodes to ensure consistency when constructing adjacency matrices. Different permutations of node indices can lead to inconsistency and higher training complexity, particularly for graph reconstruction tasks. To address this, we define a custom ordering where nodes are sorted first by event index, then by agent identifier (ID), and finally by the sequence index within the agent’s actions, ensuring a consistent representation across samples. Fig. 2 shows sample adjacency matrices with canonical and arbitrary orderings.
Figure 2: Canonical ordering (top) vs. arbitrary ordering (bottom). We sort nodes by event, agent ID, and timestep.
Node Features. Each node is annotated with a dense feature vector comprising the action, location, agent ID, normalized sequence index within the agent’s timeline, global sequence index, and Laplacian [60] positional encodings. Categorical attributes (action, location, agent ID) are mapped to learnable embeddings of dimensions $16$, 88, and $16$, respectively. The normalized time index adds a scalar feature, global index uses a 4D sinusoidal embedding, and the top-4 eigenvectors of the normalized Laplacian add structural context. Concatenating all components yields a $49$-dim node representation.
@@
Figure 3: We use two synergetic VGAEs—one for graph reconstruction and one for node feature reconstruction—whose encoders map inputs to latent spaces regularized by a shared Condition Network.
@@ 5.2 Training
We train two conditional VGAEs that jointly learn to reconstruct graph structure and node features from crowd scenario graphs; we present the model’s architecture in Fig. 3. The two models share a common encoder that produces latent embeddings, while using distinct decoders specialized for structure (VGAE-S) and features (VGAE-F). The encoder learns an approximate posterior distribution, $q(Z\mid X,C)$, where $Z_{S}$ and $Z_{F}$ are graph-level latents for structure and features, respectively. We observe that using a fixed standard normal prior, $p(Z)=\mathcal{N}(0,I)$, leads to a mismatch between $q(Z\mid X,C)$ and $p(Z)$, resulting in posterior collapse that degrades generalization. To mitigate this mismatch, we introduce a prior network that parameterizes the conditional priors $p(Z_{S}\mid C)$ and $p(Z_{F}\mid C)$, where $C$ is the condition vector. This conditional prior aligns the posterior with contextual and structural cues, enabling generation of graphs consistent with given conditions.
@@
where $\mathbf{h}_{v}^{(k)}$ is the representation of node $v$, $\mathcal{N}(v)$ its neighbors, and $e_{uv}$ the edge feature (“sequence”/“share”). After message passing, node embeddings are aggregated with global add pooling into a graph-level vector, normalized, and concatenated with the condition embedding $C$. Fully connected layers project this joint representation into the mean and log-variance of the Gaussian posterior,
$q(Z\mid X,C)=\mathcal{N}\!\big(\mu(X,C),\operatorname{diag}(\exp(\log\sigma^{2}(X,C)))\big),$ (2)
@@
Feature Decoder. The feature decoder reconstructs node features $\mathbf{X}_{G^{\prime}}$ from the latent variable $Z_{F}$, the condition vector $C$, and graph-level positional encodings. At the node level, $Z_{F}$ and $C$ are expanded and concatenated with projected positional encodings, which combine Laplacian embeddings from the share-only subgraph (capturing the connectivity in individual groups induced by “share” edges) and the full graph (capturing the overall structure), together with a normalized group identifier. The fused representation is processed by stacked GINE layers with pre-normalization and residual connections, enabling message passing over the graph.
To enforce consistency among agents that interact, node embeddings are collapsed into connected components defined by share edges, ensuring that members of the same group, at same timestep, are assigned a common action and location. Class logits are predicted at this group level and then uncollapsed back to the node level for supervision. Formally, for each node $v\in V$, the decoder outputs categorical distributions over actions and locations,
$\widehat{\mathbf{X}}^{\text{act}}_{v}=\text{softmax}(f^{\text{act}}_{\theta}(h_{v})),\qquad\widehat{\mathbf{X}}^{\text{loc}}_{v}=\text{softmax}(f^{\text{loc}}_{\theta}(h_{v})),$ (4)
where $h_{v}$ is the reconstructed embedding of node $v$ after the collapse–predict–uncollapse pipeline. These distributions are directly compared with the ground-truth labels $\mathbf{X}_{v}^{\text{act}}$ and $\mathbf{X}_{v}^{\text{loc}}$ during training.
@@
Condition Network. Both VGAE models use the same condition network, which encodes the text embedding of $S_{in}$ together with normalized counts of nodes, agents, and events, as well as an action-frequency vector, to produce the condition vector $C$. This vector is also used to train two prior networks that parameterize the conditional priors $p(Z\mid C)$ for VGAE-S and VGAE-F, yielding their respective means $\mu_{\text{prior}}$ and log-variances $\log\sigma_{\text{prior}}^{2}$.
The condition vector is integrated into both the encoding and decoding phases. In the encoder, $C$ is concatenated with the pooled graph representation to produce the posterior distribution $q(Z\mid X,C)$, parameterized by $\mu$ and $\log\sigma^{2}$. During decoding, it is concatenated with the latent variable to guide reconstruction.
@@
Training Objective. The two VGAE models are trained by maximizing the evidence lower bound (ELBO) [63], which balances reconstruction fidelity with regularization of the latent space. For the graph-structure model, VGAE-S, the reconstruction loss $\mathcal{L}_{S}$ compares the predicted adjacency $\widehat{\mathcal{A}}$ against the ground-truth adjacency $\mathcal{A}$ using a smooth L1-error loss (SmoothL1):
@@
$\mathcal{L}_{S}=\sum_{i,j}\text{SmoothL1}\!\left(\mathcal{A}_{ij},\widehat{\mathcal{A}}_{ij}\right).$ (5)
For the feature model (VGAE-F), the reconstruction loss $\mathcal{L}_{F}$ compares categorical predictions of actions and locations with ground-truth labels using cross-entropy (CELoss) at the node level:
$\mathcal{L}_{F}=\sum_{v\in V}\Big(\gamma\text{CELoss}\big(\mathbf{X}_{v}^{\text{act}},\widehat{\mathbf{X}}_{v}^{\text{act}}\big)+\text{CELoss}\big(\mathbf{X}_{v}^{\text{loc}},\widehat{\mathbf{X}}_{v}^{\text{loc}}\big)\Big),$ (6)
@@
The final training objectives are:
$\mathcal{L}_{VGAE-S}=\mathcal{L}_{S}+\beta_{S}\,\mathcal{L}_{KL_{S}},\qquad\mathcal{L}_{VGAE-F}=\mathcal{L}_{F}+\beta_{F}\,\mathcal{L}_{KL_{F}},$ (8)
where $\beta_{S}$ and $\beta_{F}$ are annealing factors that control the relative weight of the KL divergence. Both are scheduled using a cyclical annealing strategy [64], in which the KL weight is periodically ramped from zero to its maximum value over the course of training.
@@ 6 Crowd Scenario Generation
At inference time, novel crowd scenarios are generated by sampling from the learned priors using the input condition $C$. The condition includes: a textual description $S_{in}$, graph-level statistics (counts of nodes, agents, and events), and a frequency map of the available actions. From this condition, the structure latent $Z_{S}\sim p(Z_{S}\mid C)$ and the feature latent $Z_{F}\sim p(Z_{F}\mid C)$ are drawn and decoded into an adjacency matrix $\widehat{\mathcal{A}}$ and node features $\widehat{\mathbf{X}}$, respectively. This ensures that both connectivity and node attributes are contextually aligned with the full conditioning signal.
Each generated graph corresponds to a subgraph that encodes the actions and interactions of a group of agents, as introduced in Section 4.2; the framework scales naturally to arbitrary numbers of agents or groups. From the decoded graphs, we extract per-agent sequences of actions, locations, and interactions. To instantiate these scenarios in a virtual scene, the environment’s layout can either be provided manually or synthesized with a lightweight LLM query ($Q1$) (see Section 4.1). The resulting scenarios are simulated and visualized in Unity [5], with examples shown in Fig. 8 and additional animations included in the supplementary video.
@@ 7 Experiments and Evaluation
We train our model on two synthetic datasets capturing crowd scenarios in a University Campus and a Train Station. Using the bootstrapping procedure described in Section 4.1, we generate 500 scenarios per dataset. These are converted into interaction graphs as outlined in Section 4.2, yielding 22.922.9k subgraphs for University Campus and 25.125.1k for Train Station. We split each dataset into 80% training, 15% validation, and 5% testing. The code, datasets and trained models will be released upon acceptance.
For model training (cf. Fig. 3), the encoder consists of two GINE layers with hidden dimension 128, each followed by dropout ($p=0.2$). Both VGAE-S and VGAE-F use a latent dimension of 32. The graph-structure decoder is a three-layer MLP with hidden dimension 128, while the feature decoder applies two GINE layers (hidden dimension 128) followed by two linear heads that map to 13 and 8 output classes for actions and locations, respectively. The condition network projects text embeddings (768D), scalar values (3D), and action frequencies (13D) into 128 dimensions each; their concatenation is passed through an MLP to produce a 32-dimensional condition vector; the prior network maps this to the parameters of the prior distribution via two linear layers. LeakyReLU activations are used throughout, with LayerNorm applied in the encoder, both decoders, and condition network for stability. Models are trained for 500 epochs with a batch size of 128, learning rate of $5\times 10^{-4}$, linearly decayed to $1.25\times 10^{-4}$ over training, and Adam optimizer with weight decay of $3\times 10^{-4}$. The $\beta$ coefficient is annealed cyclically over 100-epoch cycles, in the ranges $[0.001,3]$ for VGAE-S and $[0.001,1]$ for VGAE-F. Training each model required approximately one hour on a single NVIDIA RTX 4070 Ti GPU.
@@ 7.1 Ablation Study
To evaluate the impact of our design choices, we compare the proposed Gen-C model against two variations of our architecture on the test set: (i) a version trained without canonical node ordering (w/o Can.), and (ii) a single VGAE model jointly trained on both structure and features (Single). Additionally, we include a constrained random baseline (Rand.), designed to support fair comparisons by generating graphs that conform to the same structural constraints as the ground truth (e.g., each node has 1–2 temporal edges). Node actions and locations are sampled from empirical priors and perturbed with mild noise, preserving marginal distributions while introducing controlled variability.
Performance is reported using the KL divergence (KLD) between the distributions of ground-truth and generated samples across both structural and semantic measures. Lower values indicate closer alignment between generated and real samples. Structural fidelity is assessed using standard graph descriptors [65]: the degree of each node $v$ is its number of incident edges; the clustering coefficient (CC) quantifies local transitivity, defined as $c(v)=\tfrac{2T(v)}{k_{v}(k_{v}-1)}$ where $T(v)$ is the number of triangles involving $v$ and $k_{v}$ its degree (with $c(v)=0$ if $k_{v}<2$); the graph diameter is the maximum shortest-path length between connected nodes; and the average path length (APL) is the mean shortest-path distance across all connected node pairs. Semantic alignment is evaluated by comparing the distributions of predicted and ground-truth action and location labels.
As shown in Table 1, Gen-C consistently achieves the lowest divergence across most metrics in both datasets. The variant without canonical ordering performs poorly on structural measures, highlighting the importance of consistent node indexing for stable learning. The single VGAE model improves over this baseline but underperforms Gen-C, especially on semantic metrics, suggesting that separating structure and feature learning yields more accurate alignment between agent interactions and their actions/locations; interestingly, the diameter statistic benefits from joint training. The random baseline, while constrained, performs poorly on graph-structure metrics but shows respectable feature-level alignment, since actions and locations are sampled from the provided input distributions. Overall, the ablation confirms that the full Gen-C framework is necessary to faithfully capture both the structural and behavioral properties of crowd scenarios.
@@
Metrics University Campus (KLD) $\downarrow$ Train Station (KLD) $\downarrow$
w/o Can. Single Rand. Gen-C w/o Can. Single Rand. Gen-C
Degree 3.496 0.062 0.595 0.056 3.289 0.051 0.468 0.033
$C$ 0.955 0.034 0.221 0.031 0.147 0.053 0.438 0.038
Diameter 3.463 0.145 0.494 0.156 3.649 0.094 0.804 0.104
APL 0.923 0.276 0.885 0.147 0.714 0.144 1.235 0.099
Action 0.180 0.358 0.225 0.177 0.128 0.340 0.368 0.106
Location 0.024 0.058 0.024 0.017 0.019 0.031 0.047 0.018
7.2 Assessment of LLM-bootstrapped Data
Gen-C leverages LLMs to synthesize crowd scenario graphs, thereby reducing the dependence on real-world labeled data. To evaluate this approach, we annotated high-level action plans of individuals interacting in Train Stations, using a curated set of YouTube videos totaling approximately 30 minutes of footage. Comparable data for university campus environments are scarce, primarily due to ethical restrictions on video collection, and existing datasets such as UCY Dataset [36] exhibit limited diversity in observed actions. Specifically, we divide each video clip into $5$-second windows and assign an action label from $\mathbf{\mathcal{A}ct}$ to each person visible in the frame.
[…]
Figure 4: Per-agent action transitions (LLM) and deviation from Real ($\Delta$) in Train Station; rows“from”, columns “to”.
As an additional statistic, we measure diversity of action plans using the normalized entropy of per-agent action sequences. Let $\mathcal{S}$ be the set of distinct action sequences and $p(s)$ their empirical frequencies. We compute $H(p)\;=\;-\sum_{s\in\mathcal{S}}p(s)\,\log p(s)$ and report the normalized value $\hat{H}=H(p)/\log|\mathcal{S}|\in[0,1]$: $\hat{H}=0$ means one pattern dominates; $\hat{H}=1$ means all patterns are equally likely. By this measure, LLM sequences closely match the annotated real-world data: $\hat{H}_{\text{LLM}}=0.949$ $(95\%\ \text{CI }[0.916,\,0.970])$ vs. $\hat{H}_{\text{Real}}=0.936$ $(95\%\ \text{CI }[0.935,\,0.939])$. The overlapping intervals indicate that LLM data follow real-world sequence variety for the given context.

### v4

@@ 4.2 Graph-Based Representation
Training a generative model for crowd scenarios requires an expressive data structure that captures agent-environment interactions and actions over time. We adopt a graph representation that meets these criteria and introduce the crowd scenario graph, a time-expanded graph modeling dynamic crowd scenarios through temporally encoded actions and interactions.
As described in Section 4.1, during simulation we collect $Rec_{i}$ for each agent $i$. Iterating through $Rec_{i}$, we create a node $V_{i}^{t}=(i,A_{i}^{t},L_{i}^{t})$, where $t$ is the current timestep, $A_{i}^{t}\in\mathrm{A}ct$, and $L_{i}^{t}\in\mathrm{L}oc$; the combination of $A_{i}^{t}$ and $L_{i}^{t}$ defines an agent-environment interaction at $t$. We set the action sequence of $i$ using a “sequence edge” $E_{i}^{t}=(V_{i}^{t},V_{i}^{t-1})$. Next, agent-agent interactions are captured by connecting nodes from different agents that share an action at a given $t$. A “share edge” $E_{i,j}^{t}=(V_{i}^{t},V_{j}^{t})$ is created for each pair of agents $i$ and $j$ interacting at $t$. At this stage, nodes of each agent $i$ are sequentially connected into a chain. If the chain has no “shared edges,” it forms an individual subgraph $H_{k}$ corresponding to agent $i$. When at least one “shared edge” exists, the chains of multiple agents merge into a subgraph representing a group of agents. Thus, for each $CS$, we construct an undirected graph $G=(V,E)$, consisting of multiple subgraphs $\mathcal{H}=\{H_{1},H_{2},\dots\}$, where $V$ is the set of nodes and $E$ the set of edges. Finally, each subgraph $H_{k}\in\mathcal{H}$ is treated as an individual training sample $(S_{in},H_{k},\text{agents}_{k})$, where $H_{k}$ is the subgraph for group $k$, and $\text{agents}_{k}$ the agents in that group. Figure 4 presents a snapshot of a subgraph with five agents from a crowd scenario graph.
@@ 5.1 Data Preprocessing
Node Ordering. We enforce a canonical ordering of nodes to ensure consistency when constructing adjacency matrices. Different permutations of node indices can lead to inconsistency and higher training complexity, particularly for graph reconstruction tasks. To address this, we define a custom ordering where nodes are sorted first by event index, then by agent identifier (ID), and finally by the sequence index within the agent’s actions, ensuring a consistent representation across samples. Figure 4 shows sample adjacency matrices with canonical and arbitrary orderings. Note that this ordering standardizes node indexing for stable training but does not directly enforce logical action sequences, which are instead captured implicitly through the LLM-bootstrapped training data.
Node Features. Each node is represented by a dense feature vector encoding action, location, agent ID, normalized position in the agent timeline, global sequence index, and Laplacian [6] positional encodings. Categorical features are mapped to learnable embeddings of size $16$, 88, and $16$, respectively. The normalized time index contributes one scalar, the global index a 4D sinusoidal embedding, and the top-4 eigenvectors of the normalized Laplacian provide structural context, yielding a $49$-dimensional node representation.
@@
Figure 5: Our architecture comprises two synergistic VGAEs: one for graph reconstruction and one for node feature reconstruction, with encoders mapped to latent spaces regularized by a shared Condition Network.
@@ 5.2 Training
We train two conditional VGAEs that jointly learn to reconstruct graph structure and node features from crowd scenario graphs; we present the model’s architecture in Figure 5. The two models share a common encoder that produces latent embeddings, while using distinct decoders specialized for structure (VGAE-S) and features (VGAE-F). The encoder learns an approximate posterior distribution, $q(Z\mid X,C)$, where $Z_{S}$ and $Z_{F}$ are graph-level latents for structure and features, respectively. We observe that using a fixed standard normal prior, $p(Z)=\mathcal{N}(0,I)$, leads to a mismatch between $q(Z\mid X,C)$ and $p(Z)$, resulting in posterior collapse that degrades generalization. To mitigate this mismatch, we introduce a prior network that parameterizes the conditional priors $p(Z_{S}\mid C)$ and $p(Z_{F}\mid C)$, where $C$ is the condition vector. This conditional prior aligns the posterior with contextual and structural cues, enabling generation of graphs consistent with given conditions.
@@
where $\mathbf{h}_{v}^{(k)}$ is the representation of node $v$, $\mathcal{N}(v)$ its neighbors, and $e_{uv}$ the edge feature (“sequence”/“share”). After message passing, node embeddings are aggregated with global add pooling into a single vector, normalized, and concatenated with the condition embedding $C$. Fully connected layers project this joint representation into the mean and log-variance of the Gaussian posterior,
$q(Z\mid X,C)=\mathcal{N}\Big(\mu(X,C),\operatorname{diag}\big(\exp(\log\sigma^{2}(X,C))\big)\Big),$ (2)
@@
Feature Decoder. The feature decoder reconstructs node features $\mathbf{X}_{G^{\prime}}$ from the latent variable $Z_{F}$, the condition vector $C$, and graph-level positional encodings. At the node level, $Z_{F}$ and $C$ are expanded and concatenated with projected positional encodings, which combine Laplacian embeddings from the share-only subgraph (capturing the connectivity in individual groups induced by “share” edges) and the full graph (capturing the overall structure), together with a normalized group identifier. The fused representation is processed by stacked GINE layers with pre-normalization and residual connections, enabling message passing over the graph. To enforce consistency among agents that interact, node embeddings are collapsed into connected components defined by share edges, ensuring that members of the same group, at same timestep, are assigned a common action and location. Class logits are predicted at this group level and then uncollapsed back to the node level for supervision. Formally, for each node $v\in V$, the decoder outputs categorical distributions over actions and locations as: $\widehat{\mathbf{X}}^{\text{act}}_{v}=\text{softmax}(f^{\text{act}}_{\theta}(h_{v}))$, and $\widehat{\mathbf{X}}^{\text{loc}}_{v}=\text{softmax}(f^{\text{loc}}_{\theta}(h_{v}))$, where $h_{v}$ is the reconstructed embedding of node $v$ after the collapse–predict–uncollapse pipeline. These distributions are directly compared with the ground-truth labels $\mathbf{X}_{v}^{\text{act}}$ and $\mathbf{X}_{v}^{\text{loc}}$ during training.
@@
Condition Network. Both VGAE models use the same condition network, which encodes the text embedding of $S_{in}$ together with normalized counts of nodes, agents, and events, as well as an action-frequency vector, to produce the condition vector $C$. This vector is also used to train two prior networks that parameterize the conditional priors $p(Z\mid C)$ for VGAE-S and VGAE-F, yielding their respective means $\mu_{\text{prior}}$ and log-variances $\log\sigma_{\text{prior}}^{2}$. The condition vector is integrated into both the encoding and decoding phases. In the encoder, $C$ is concatenated with the pooled graph representation to produce the posterior distribution $q(Z\mid X,C)$, parameterized by $\mu$ and $\log\sigma^{2}$. During decoding, it is concatenated with the latent variable to guide reconstruction.
@@
Training Objective. VGAE models maximize the ELBO [22], balancing reconstruction fidelity with latent space regularization. VGAE-S uses smooth L1-error (SmoothL1) loss $\mathcal{L}_{S}$ to compare predicted adjacency $\widehat{\mathcal{A}}$ against ground-truth $\mathcal{A}$, while VGAE-F applies node-level cross-entropy (CELoss) $\mathcal{L}_{F}$ for action and location predictions. The losses are defined as:
@@
$\begin{split}\mathcal{L}_{S}&=\sum_{i,j}\text{SmoothL1}\!\left(\mathcal{A}_{ij},\widehat{\mathcal{A}}_{ij}\right),\\ \mathcal{L}_{F}&=\sum_{v\in V}\Big(\gamma\text{CELoss}\big(\mathbf{X}_{v}^{\text{act}},\widehat{\mathbf{X}}_{v}^{\text{act}}\big)+\text{CELoss}\big(\mathbf{X}_{v}^{\text{loc}},\widehat{\mathbf{X}}_{v}^{\text{loc}}\big)\Big).\end{split}$ (4)
@@
The final training objectives are: $\mathcal{L}_{VGAE-S}=\mathcal{L}_{S}+\beta_{S}\,\mathcal{L}_{KL_{S}}$ and $\mathcal{L}_{VGAE-F}=\mathcal{L}_{F}+\beta_{F}\,\mathcal{L}_{KL_{F}}$, where $\beta_{S}$ and $\beta_{F}$ are annealing factors that control the relative weight of the KL divergence.
@@ 6 Crowd Scenario Generation
At inference time, novel crowd scenarios are generated by sampling from the learned priors using the input condition $C$. The condition includes: a textual description $S_{in}$, graph-level statistics (counts of nodes, agents, and events), and a frequency map of the available actions. From this condition, the structure latent $Z_{S}\sim p(Z_{S}\mid C)$ and the feature latent $Z_{F}\sim p(Z_{F}\mid C)$ are drawn and decoded into an adjacency matrix $\widehat{\mathcal{A}}$ and node features $\widehat{\mathbf{X}}$, respectively. This ensures that both connectivity and node attributes are contextually aligned with the full conditioning signal. Each generated graph corresponds to a subgraph that encodes the actions and interactions of a group of agents, as introduced in Section 4.2; the framework scales naturally to arbitrary numbers of agents or groups. From the decoded graphs, we extract per-agent sequences of actions, locations, and interactions. To instantiate these scenarios in a virtual scene, the environment’s layout can either be provided manually or synthesized with a lightweight LLM query ($Q1$) (see Section 4.1).
@@ 7 Experiments and Evaluation
We train on two synthetic crowd datasets, University Campus and Train Station. Following the procedure described in Section 4.1, we generate 500 scenarios per dataset and convert them into graphs as in Section 4.2, obtaining 22.922.9K subgraphs for University Campus and 25.125.1K for Train Station. Each dataset is split into 80% training, 15% validation, and 5% testing. The source code, datasets, and trained models can be accessed at https://github.com/apanay20/Gen-C.
For training (see Figure 5), the encoder uses two GINE layers with hidden size 128, each followed by dropout ($p=0.2$). Both VGAE-S and VGAE-F use a 32-dimensional latent space. The graph decoder is a three-layer MLP with hidden size 128, while the feature decoder uses two GINE layers (hidden size 128) followed by two linear heads predicting 13 action classes and 8 location classes. The condition network projects text embeddings (768D), scalar inputs (3D), and action frequencies (13D) to 128 dimensions each; their concatenation is processed by an MLP to produce a 32-dimensional condition vector, which the prior network maps to prior distribution parameters through two linear layers. LeakyReLU is used throughout, and LayerNorm is applied in the encoder, both decoders, and condition network for stability. Models are trained for 500 epochs with batch size 128, learning rate $5\times 10^{-4}$ linearly decayed to $1.25\times 10^{-4}$, and Adam with weight decay $3\times 10^{-4}$. The $\beta_{S}$ and $\beta_{F}$ coefficients are cyclically annealed over 100-epoch cycles in the ranges $[0.001,3]$ and $[0.001,1]$, respectively. Training each model takes about one hour on a single NVIDIA RTX 4070 Ti GPU.
@@ 7.1 Ablation Study
To evaluate the impact of our design choices, we compare the proposed Gen-C model against two variations of our architecture on the test set: (i) a version trained without canonical node ordering (w/o Canonical), and (ii) a single VGAE model jointly trained on both structure and features (Single). Additionally, we include a constrained random baseline (Random), designed to support fair comparisons by generating graphs that conform to the same structural constraints as the ground truth (e.g., each node has 1-2 temporal edges). Node actions and locations are sampled from empirical priors and perturbed with mild noise, preserving marginal distributions while introducing controlled variability. Performance is reported using the KL divergence (KLD) between the distributions of ground-truth and generated samples across both structural and semantic measures. Lower values indicate closer alignment between generated and real samples. Structural fidelity is assessed using standard graph descriptors [40]: the degree of each node $v$ is its number of incident edges; the clustering coefficient quantifies local transitivity, defined as $c(v)=\tfrac{2T(v)}{k_{v}(k_{v}-1)}$ where $T(v)$ is the number of triangles involving $v$ and $k_{v}$ its degree (with $c(v)=0$ if $k_{v}<2$); the graph diameter is the maximum shortest-path length between connected nodes; and the average path length is the mean shortest-path distance across all connected node pairs. Semantic alignment is evaluated by comparing the distributions of predicted and ground-truth action and location labels.
As shown in Table 1, Gen-C consistently achieves the lowest divergence across most metrics in both datasets. The variant without canonical ordering performs poorly on structural measures, highlighting the importance of consistent node indexing for stable learning. The single VGAE model improves over this baseline but underperforms Gen-C, especially on semantic metrics, suggesting that separating structure and feature learning yields more accurate alignment between agent interactions and their actions/locations; interestingly, the diameter statistic benefits from joint training. The random baseline, while constrained, performs poorly on graph-structure metrics but shows respectable feature-level alignment, since actions and locations are sampled from the input distributions. Overall, the ablation confirms that the full framework is necessary to faithfully capture both structural and behavioral properties of crowd scenarios.
@@
Metric University Campus (KLD) $\downarrow$ Train Station (KLD) $\downarrow$
w/o Canonical Single Random Gen-C w/o Canonical Single Random Gen-C
Degree 3.4955 0.0616 0.5953 0.0560 3.2891 0.0510 0.4680 0.0332
Clustering Coefficient 0.9550 0.0341 0.2212 0.0308 0.1472 0.0529 0.4381 0.0384
Diameter 3.4630 0.1448 0.4942 0.1559 3.6493 0.0937 0.8044 0.1038
Average Path Length 0.9234 0.2761 0.8854 0.1466 0.7142 0.1436 1.2349 0.0990
Action 0.1800 0.3579 0.2251 0.1773 0.1280 0.3399 0.3681 0.1057
Location 0.0242 0.0576 0.0244 0.0171 0.0188 0.0314 0.0465 0.0177
To further assess the behavioral fidelity of Gen-C, we conduct a random perturbation experiment in which nodes and edges are randomly added or removed from ground-truth graphs following the proposed canonical ordering (Section 5.1); features are randomly sampled from a uniform distribution. We use four semantic-oriented metrics capturing temporal and relational dynamics: (i) action transition KLD and location transition KLD (sequential consistency); (ii) joint action-location KLD (co-occurrence distributions between actions and locations); and (iii) interaction participation rate (proportion of agents engaged in at least one social interaction). As shown in Table 2, Gen-C consistently outperforms the random baseline (Perturb) across both datasets. The model achieves significantly lower KLD scores and maintains interaction participation rates closely aligned with the ground-truth (GT). These findings highlight the value of a trained model over random perturbations, showing that Gen-C effectively captures the semantic regularities underlying agent behaviors.
Table 2: Random perturbation comparison.
Metric University Campus Train Station
Perturb Gen-C GT Perturb Gen-C GT
Act. Trans. $\downarrow$ 0.584 0.250 - 0.679 0.239 -
Loc. Trans. $\downarrow$ 0.566 0.135 - 0.767 0.427 -
Joint KLD $\downarrow$ 0.899 0.248 - 1.121 0.391 -
Interact Rate 0.176 0.667 0.699 0.168 0.627 0.662